OpenStudy (baozipotato97):

Identify the point (x1,y1) from the equation: y+8=3(x-2) A. (-2,8) B. (8,-2) C. (-8, 2) D. (2,-8) Thank you!

OpenStudy (kidrah69):

This is the point slope equation \[\huge y-y_1=m(x-x_1)\] where m=slope Use it to match up the numbers to the corresponding letter :-) Also, Welcome to Opnestudy!!

OpenStudy (kidrah69):

Keep in mind that a negative x negative= positive

OpenStudy (baozipotato97):

Thank you~! :)

OpenStudy (kidrah69):

yw! :D